# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### v1

@@ Abstract
We present SciCoQA, a dataset for detecting discrepancies between scientific publications and their codebases to ensure faithful implementations. We construct SciCoQA from GitHub issues and reproducibility papers, and to scale our dataset, we propose a synthetic data generation method for constructing paper-code discrepancies. We analyze the paper-code discrepancies in detail and propose discrepancy types and categories to better understand the occurring mismatches. In total, our dataset consists of 611 paper-code discrepancies (81 real, 530 synthetic), spanning diverse computational science disciplines, including AI, Physics, Quantitative Biology, and others. Our evaluation of 21 LLMs highlights the difficulty of SciCoQA, particularly for instances involving omitted paper details, long-context inputs, and data outside the models’ pre-training corpus. The best performing model in our evaluation, GPT-5, can only detect $45.7\%$ of real-world paper-code discrepancies.
@@
Tim Baumgärtner  and  Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE [Uncaptioned image] Code: https://github.com/ukplab/scicoqa [Uncaptioned image] Data: https://hf.co/datasets/ukplab/scicoqa
Data also available at: https://tudatalib.ulb.tu-darmstadt.de/handle/tudatalib/4994
@@ 1 Introduction
However, the availability of code does not guarantee reproducibility nor consistency with the scientific text (as showcased in Fig. 1). In practice, implementation details can diverge from their descriptions, introducing performance variations that go unreported (Henderson et al., 2018). These discrepancies manifest in troubling ways: from “mathiness,” where equations simulate technical depth while actual gains stem from undocumented tricks (Lipton and Steinhardt, 2019), to evaluation metrics that differ in implementation (Post, 2018), rendering scientific comparisons invalid.
@@
Furthermore, the reliance on manual review is becoming increasingly impractical as science begins to scale via agentic systems which develop ideas, generate and execute code, and produce scientific articles autonomously (Lu et al., 2024; Tang et al., 2025; Weng et al., 2025b), a trend recently validated by peer-reviewed acceptance to an ICLR workshop (Yamada et al., 2025). While this direction has the potential to accelerate scientific discovery, it makes oversight increasingly challenging as humans cannot verify the rapidly expanding volume of output (Bowman et al., 2022). Moreover, the faithfulness of these systems is not guaranteed; LLM agents suffer from limited context (Liu et al., 2024b), experience compounding errors (Dziri et al., 2023; Mirzadeh et al., 2025), and struggle with independent self-correction (Tyen et al., 2024; Wu et al., 2024).
Consequently, ensuring the reliability of science at this scale demands automated tools capable of verifying that the code faithfully implements the methods reported in the scientific paper. To this end, we introduce the SciCoQA dataset for a novel, real-world quality assurance task: detecting discrepancies between scientific papers and code in the computational sciences. Benchmarking LLMs on this task allows us to measure their potential for automated quality assurance workflows in science. To construct our dataset, we leverage issues reported in code repositories and papers dedicated to reproducing existing work, such as those from reproducibility challenges and conference tracks. While these sources are realistic and of high quality, they are sparse and limited to CS and AI. Therefore, we scale the data to various computational science domains (e.g., Physics and Quantitative Biology) by generating modifications in codebases, thereby creating discrepancies between the paper and its implementation. Finally, we benchmark open-weight and proprietary LLMs to determine whether current systems can be deployed for detecting paper-code discrepancies. Our evaluation reveals that the best-performing models are precise in identifying discrepancies; however, their recall is too low to guarantee that the code faithfully implements the paper, highlighting the need to develop more effective models and systems for this critical task.
@@ 2 Related Work
Recently, LLMs have been employed to detect errors in scientific papers. In an early work, Liu and Shah (2023) created short CS papers with logical errors and prompted LLMs to identify the issues. They found that while models could produce fluent reviews, they struggled to identify these fundamental errors without explicit instructions in the prompt. However, their reliance on manually crafted papers limits scalability. Therefore, Dycke and Gurevych (2026) and Xi et al. (2025) propose dedicated pipelines that generate modifications to papers that invalidate their key claims and evaluate whether the soundness score of generated reviews decreases or whether LLMs can spot these errors. Both find that LLMs struggle to detect most introduced errors. Besides constructing erroneous papers, recent research has also focused on identifying real-world issues. Specifically, Son et al. (2025) and Zhang and Abernethy (2025) use the authors’ retraction notes from WithdrarXiv (Rao et al., 2024) and PubPeer to obtain data of errors in scientific publication. While realistic, these retraction notes are often unspecific or describe errors on a high level only (e.g., "Theorem 2 is incorrect"). Recently, Bianchi et al. (2025) developed a system to detect inconsistencies within the text of AI papers (e.g., incorrect calculations in tables, errors in equations, and imprecise definitions) and then manually verify the detections. They found their system to be relatively precise, but struggled with recall, i.e., detecting all issues in a paper. They also noted an increasing trend of errors in scientific publications over the years, highlighting the importance of quality assurance in science.
@@
Figure 2: Overview of the data collection process of SciCoQA. We source real-world data from Reproducibility papers and GitHub issues. For the former, paper-code discrepancies are extracted from the paper with GPT-5, for the latter, issues are pre-filtered using Qwen3. Next, all candidates are manually filtered to remove any that do not fit our discrepancy definition. Finally, all paper-code discrepancies are verified with GPT-5. For synthetic data, we generate discrepancies using GPT-5 for AI and other computational domains.
@@
While these works share a similar spirit to ours, i.e., identifying errors in science, they are limited by solely analyzing the paper. SciCoQA extends beyond the scientific text and considers the code as well, enabling the detection of cross-modal error types, such as implementations that deviate from the paper description or omissions in the paper or code that are crucial for understanding and reproducibility. Furthermore, SciCoQA utilizes constructed errors to scale up the quantity and diversity in the computational science domain, while also providing challenging real-world discrepancies between paper and code.
Comment-Code Inconsistency
@@
Our task is related to Code-Comment Inconsistency (CCI) detection, where natural language comments that no longer match a piece of code, e.g., after adding a feature to a function, need to be identified. Early CCI research relied on rule-based systems to identify outdated comments (Ratol and Robillard, 2017). Subsequently, bi-encoders (Rabbi and Siddik, 2020) and cross-encoders (Steiner and Zhang, 2022) have been employed. Panthaplackel et al. (2021) further extended this to dynamic settings, learning to detect inconsistency arising from code updates. Recently, LLMs have been deployed for both detection and rectification (Dau et al., 2024; Rong et al., 2025). These methods deal with local and small inputs, typically taking a single function and an inline comment or docstring as input to detect the CCI. SciCoQA is crucially much more ambitious: it requires performing a global alignment, by reasoning over a dense scientific paper and a long, multi-file code repository to find inconsistencies.
@@
Automation in science is rapidly accelerating. Early work targeted specific components of the research cycle, including ideation (Si et al., 2025; Baek et al., 2025), literature review Wang et al. (2024); Liang et al. (2025), coding (Tian et al., 2024; Gandhi et al., 2025), and writing (Liang et al., 2024; Weng et al., 2025a). A recent trend involves generating entire code repositories from a paper (Starace et al., 2025; Seo et al., 2025; Xiang et al., 2025) or automating the scientific method end-to-end, giving rise to "AI Scientists." However, for both these tasks, validating whether the generated code faithfully reflects the paper is challenging. For example, PaperBench (Starace et al., 2025) relies on high-quality, but expensive, manual rubrics specific to each paper to verify implementation nuances. Others use generic LLM-judges without ground truth validation Tang et al. (2025), evaluate on whether the code runs (Weng et al., 2025b; Xiang et al., 2025), or assign reproducibility scores based on execution outcomes Siegel et al. (2024); Hu et al. (2025); Ye et al. (2025). Critically, these metrics can be misleading; a generated codebase may execute perfectly and achieve high performance while implementing a method that differs fundamentally from the paper’s scientific description (Beel et al., 2025). SciCoQA addresses this gap by providing a ground truth dataset of paper-code discrepancies, enabling the development of robust quality assurance models that can verify the faithfulness of (generated) papers to (generated) code.
@@ 3 SciCoQA
We define a paper-code discrepancy as a semantic conflict between the scientific method described in the publication and its actual implementation in the codebase, such that the code does not faithfully reproduce the reported method. This mismatch must be meaningful, implying a fundamental alteration to the scientific logic, experimental protocol, or mathematical formulation described in the text. These discrepancies manifest as three distinct types: differences, where the code implements a logic distinct from the paper’s description (e.g., L1 vs. L2 normalization) paper omissions, where the code includes critical components missing from the text, or code omissions, where a step described in the paper is absent from the repository. We distinguish these discrepancies from engineering artifacts: We exclude bugs, as these are independent of the paper’s scientific description. Similarly, mismatches in default hyperparameters are not considered discrepancies if the code supports the paper’s settings via configuration files or CLI arguments. Finally, we exclude trivial implementation details that are standard engineering practices typically omitted from scientific descriptions (e.g., adding noise to a denominator for numerical stability).
The paper-code discrepancies, according to our definition, can occur from distinct authorship (where the paper writer differs from the engineer implementing the code), simplifications made in the text for readability, or code updates, such as experimental variations, that were not propagated back to the manuscript.
@@ 3.1 Data Collection
We first identify repositories that reference a research paper in their homepage or description field on GitHub, restricted to projects published between 2020 and 2025 and repositories with at least one issue (see §A for details). [^1] The crawled repositories are not always the official implementation by the authors, but sometimes also reproductions. They have been published, for example, when there is no official implementation or the reimplementation uses a different framework. We still consider these repositories as researchers might choose them to base their experiments on. For each, we crawl all associated issues, yielding $1,890$ repositories with a total of $10,636$ issues. To process this large volume efficiently, we automatically classify issues with Qwen3 4B Thinking (Yang et al., 2025), prompting the model to determine whether an issue reports a paper–code discrepancy or not, resulting in $232$ candidates (prompt in §B.1). All candidates are subsequently manually filtered to ensure they meet our definition of a paper-code discrepancy, yielding $59$ discrepancies.
@@
We additionally collect reproducibility reports from the ML Reproducibility Challenge (Pineau et al., 2019; Sinha et al., 2020, 2022, 2023) which invites participants to reproduce papers from leading ML, CV, and NLP conferences (e.g., NeurIPS, ICML, ICLR, CVPR, ICCV, EMNLP, ACL). We also include papers from the reproducibility tracks of SIGIR and ECIR from 2020–2025. We retain only reports that (1) reproduce a single, open-access paper, and (2) reproduce papers with an open-source implementation, resulting in $171$ reproducibility papers. From each report, we extract mentions of paper–code discrepancies using GPT-5 resulting in $132$ candidates (prompt in §B.3). Each extracted discrepancy is manually verified to confirm adherence to our definition, of which $65$ withstand our filter.
@@
Finally, all manually filtered discrepancies are validated by GPT-5, which checks the discrepancy’s existence given the original paper, codebase, and the GitHub issue or the description extracted from the reproducibility paper (prompt in §B.2 and §B.4). While the manual verification step ensures adherence to the discrepancy definition, this step acts as a high-precision filter. By retaining only discrepancies that can be verified given the explicit evidence, we prioritize the precision and objectivity of the ground truth. In total, we obtain $81$ discrepancies from $67$ papers, where $37$ discrepancies originate from GitHub issues and $44$ from reproducibility papers. During this final verification step, GPT-5 is also tasked with generating a standardized description of discrepancies. We prompt the model to output 3-8 sentences and state (1) what the paper describes, (2) what the code implements, and (3) where the difference is. We use these generated descriptions as the ground truth to ensure the same format and level of verbosity for paper-code discrepancies.
@@
To scale the data in size and beyond the CS/AI domain, we employ synthetic data generation. Specifically, based on our previous GitHub crawl, we sample repositories linked to an arXiv paper and those with permissive code licenses (i.e., MIT, Apache 2.0, CC-BY, BSD), as we will be redistributing their code partially. We randomly select $102$ repositories where the paper’s arXiv subject is CS (balanced over Machine Learning, cs.LG, Computer Vision, cs.CV, and Natural Language Processing, cs.CL), and $102$ for non-CS papers. [^2] We manually checked the repositories and removed codebases from our sample that were not suitable, such as repositories for survey papers or with very few files. Next, we prompt GPT-5 with the paper and code to generate five code diffs according to our discrepancy definition (prompt in §B.5). We then sample up to three of these changes, ensuring that they do not manipulate the same file and that the generated diff can be found by exact match in the original code. This yields a total of $530$ discrepancies across the $204$ papers, of which $280$ discrepancies are from the CS domain, and $250$ are from other computational domains, such as Electrical Engineering and Systems Science ($112$ discrepancies), Physics ($49$), or Statistics ($39$).
Example data, including discrepancy type and category annotations, are shown in Table 7. Examples from the synthetic subset, including the code changes, are shown in Table 8. We release the SciCoQA dataset publicly under $C$-BY-4.0.

### v2

@@ Abstract
We present SciCoQA, a dataset for detecting discrepancies between scientific publications and their codebases to ensure faithful implementations. We construct SciCoQA from GitHub issues and reproducibility papers, and to scale our dataset, we propose a synthetic data generation method for constructing paper-code discrepancies. We analyze the paper-code discrepancies in detail and propose discrepancy types and categories to better understand the occurring mismatches. In total, our dataset consists of $635$ paper-code discrepancies ($92$ real, $543$ synthetic), covering the AI domain from real-world data and extending to Physics, Quantitative Biology, and other computational sciences through synthetic data. Our evaluation of $22$ LLMs demonstrates the difficulty of SciCoQA, particularly for instances involving omitted paper details, long-context inputs, and data outside the models’ pre-training corpus. The best-performing models in our evaluation, Gemini 3.1 Pro and GPT-5 Mini, detect only $46.7\%$ of real-world paper-code discrepancies.
@@
Tim Baumgärtner  and  Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science, TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE [Uncaptioned image] Code: https://github.com/ukplab/scicoqa [Uncaptioned image] Data: https://hf.co/datasets/ukplab/scicoqa [Uncaptioned image] Blog: https://ukplab.github.io/scicoqa
@@ 1 Introduction
However, the availability of code does not guarantee reproducibility nor consistency with the scientific text (as showcased in Fig. 1). In practice, implementation details can diverge from their descriptions, introducing performance variations that go unreported (Henderson et al., 2018). These discrepancies manifest in troubling ways: from “mathiness,” where equations simulate technical depth while actual gains stem from undocumented tricks (Lipton and Steinhardt, 2019), to evaluation metrics that differ in implementation (Post, 2018; Marie et al., 2021), rendering scientific comparisons invalid.
@@
Furthermore, the reliance on manual review is becoming increasingly impractical as science begins to scale via agentic systems which develop ideas, generate and execute code, and produce scientific articles autonomously (Boiko et al., 2023; Lu et al., 2024; Tang et al., 2025; Weng et al., 2025b), a trend recently validated by peer-reviewed acceptance to an ICLR workshop (Yamada et al., 2025). While this direction has the potential to accelerate scientific discovery, it makes oversight increasingly challenging as humans cannot verify the rapidly expanding volume of output (Bowman et al., 2022). Moreover, the faithfulness of these systems is not guaranteed as LLMs suffer from limited context (Liu et al., 2024b), experience compounding errors (Dziri et al., 2023; Mirzadeh et al., 2025), and struggle with independent self-correction (Tyen et al., 2024; Wu et al., 2024).
Consequently, ensuring the reliability of science at this scale demands automated tools capable of verifying that the code faithfully implements the methods reported in the scientific paper. To this end, we introduce the SciCoQA dataset for a novel, real-world quality assurance task: detecting discrepancies between scientific papers and code in the computational sciences. Benchmarking LLMs on this task allows us to measure their potential for automated quality assurance workflows in science. To construct our dataset, we leverage issues reported in code repositories and papers dedicated to reproducing existing work, such as those from reproducibility challenges and conference tracks. While these sources are realistic and of high quality, they are sparse and limited to CS and AI. Therefore, we scale the data to various computational science domains (e.g., Physics and Quantitative Biology) by generating modifications in codebases, creating discrepancies between the paper and its implementation. Finally, we benchmark open-weight and proprietary LLMs to determine whether current systems can be deployed for detecting paper-code discrepancies. Our evaluation reveals that the best-performing models are precise in identifying discrepancies; however, their recall is too low to guarantee that the code faithfully implements the paper, highlighting the need to develop more effective models and systems for this critical task.
@@ 2 Related Work
Recently, LLMs have been employed to detect errors in scientific papers. In an early work, Liu and Shah (2023) created short CS papers with logical errors and prompted LLMs to identify the issues. They found that while models could produce fluent reviews, they struggled to identify these fundamental errors without explicit instructions in the prompt. However, their reliance on manually crafted papers limits scalability. Therefore, Dycke and Gurevych (2025) and Xi et al. (2025) propose dedicated pipelines that generate modifications to papers that invalidate their key claims and evaluate whether the soundness score of generated reviews decreases or whether LLMs can spot these errors. Both find that LLMs struggle to detect most introduced errors. Besides constructing erroneous papers, recent research has also focused on identifying real-world issues. Specifically, Son et al. (2025) and Zhang and Abernethy (2025) use the authors’ retraction notes from WithdrarXiv (Rao et al., 2024) and PubPeer to obtain data of errors in scientific publication. While realistic, these retraction notes are often unspecific or describe errors on a high level (e.g., “Theorem 2 is incorrect”). Recently, Bianchi et al. (2025) developed a system to detect inconsistencies within the text of AI papers (e.g., incorrect calculations in tables, errors in equations, and imprecise definitions) and then manually verify the detections. They found their system to be relatively precise, but struggled with recall, i.e., detecting all issues in a paper. They also noted an increasing trend of errors in scientific publications over the years, highlighting the importance of quality assurance in science.
@@
Figure 2: Overview of the data collection process of SciCoQA. We source real-world data from reproducibility papers and GitHub issues. For the former, paper-code discrepancies are extracted from the paper with GPT-5, for the latter, issues are pre-filtered using Qwen3. Next, all candidates are manually filtered to remove any that do not fit our discrepancy definition. Finally, all paper-code discrepancies are verified with Gemini 3.1 and GPT-5. For synthetic data, we generate discrepancies using GPT-5 for AI and other computational domains.
@@
While these works share a similar spirit to ours, i.e., identifying errors in science, they are limited by solely analyzing the paper. SciCoQA extends beyond the scientific text and considers the code as well, enabling the detection of cross-modal error types, such as implementations that deviate from the paper description or omissions in the paper or code that are crucial for understanding and reproducibility. Furthermore, SciCoQA utilizes constructed errors to scale the quantity and diversity in the computational science domain, while also providing challenging real-world discrepancies between paper and code.
Code-Comment Inconsistency
@@
Our task is related to Code-Comment Inconsistency (CCI) detection, where natural language comments that no longer match a piece of code need to be identified, e.g., after adding a feature to a function. Early CCI research relied on rule-based systems to identify outdated comments (Ratol and Robillard, 2017). Subsequently, bi-encoders and cross-encoders have been employed (Rabbi and Siddik, 2020; Steiner and Zhang, 2022). Panthaplackel et al. (2021) further extended CCI to dynamic settings, learning to detect inconsistency arising from code updates. Recently, LLMs have been deployed for both detection and rectification (Dau et al., 2024; Rong et al., 2025). These methods deal with local and small inputs, typically taking a single function and an inline comment or docstring as input to detect the CCI. SciCoQA is crucially much more ambitious: it requires performing a global alignment, by reasoning over a dense scientific paper and a long, multi-file code repository to find inconsistencies.
@@
Automation in science is rapidly accelerating (Zheng et al., 2025). Early work targeted specific components of the research cycle, including ideation (Si et al., 2025; Baek et al., 2025), literature review Wang et al. (2024); Liang et al. (2025), coding (Tian et al., 2024; Gandhi et al., 2025), and writing (Liang et al., 2024; Weng et al., 2025a). A recent trend involves generating entire code repositories from a paper (Starace et al., 2025; Seo et al., 2025; Xiang et al., 2025) or automating the scientific method end-to-end, giving rise to “AI Scientists.” However, for both these tasks, validating whether the generated code faithfully reflects the paper is challenging. For example, PaperBench (Starace et al., 2025) relies on high-quality, but expensive, manual rubrics specific to each paper to verify implementation nuances. Others use generic LLM-judges without ground truth validation Tang et al. (2025), evaluate on whether the code runs (Weng et al., 2025b; Xiang et al., 2025), or assign reproducibility scores based on execution outcomes Siegel et al. (2024); Hu et al. (2025); Ye et al. (2025). Critically, these metrics can be misleading; a generated codebase may execute perfectly and achieve high performance while implementing a method that differs fundamentally from the paper’s scientific description (Beel et al., 2025). SciCoQA addresses this gap by providing a ground truth dataset of paper-code discrepancies, enabling the development of robust quality assurance models that can verify the faithfulness of (generated) papers to (generated) code. Thus, SciCoQA takes a step toward verification-oriented and reproducibility tooling essential for sustaining peer review integrity and science at scale (Wei et al., 2025b; Woodruff et al., 2026; You et al., 2026).
@@ 3 SciCoQA
We define a paper-code discrepancy as a semantic conflict between the scientific method described in the publication and its actual implementation in the codebase, such that the code does not faithfully reproduce the reported method. This mismatch must be meaningful, implying a fundamental alteration to the scientific logic, experimental protocol, or mathematical formulation described in the text. These discrepancies manifest as three distinct types: differences, where the code implements a logic distinct from the paper’s description (e.g., L1 vs. L2 normalization), paper omissions, where the code includes critical components missing from the text, or code omissions, where a step described in the paper is absent from the repository. We distinguish these discrepancies from engineering artifacts: We exclude bugs, as these are independent of the paper’s scientific description. Similarly, mismatches in default hyperparameters are not considered discrepancies if the code supports the paper’s settings via configuration files or CLI arguments. Finally, we exclude trivial implementation details that are standard engineering practices typically omitted from scientific descriptions (e.g., adding noise to a denominator for numerical stability).
The paper-code discrepancies, according to our definition, can occur from distinct authorship (where the paper writer differs from the engineer implementing the code), simplifications made in the text for readability, or code updates or experiments, that were not propagated to the manuscript.
@@ 3.1 Data Collection
We first identify repositories that reference a research paper in their homepage or description field on GitHub, restricted to projects published between 2020 and 2025 and repositories with at least one issue (see §A.1 for details). [^1] The crawled repositories are not always the official implementation by the authors, but sometimes reproductions. They were published, for example, when there is no official implementation or the reimplementation uses a different framework. We still consider these repositories as researchers might use them to base their experiments on. For each, we crawl all associated issues, yielding 1,8901{,}890 repositories with a total of 10,63610{,}636 issues. To process this large volume efficiently, we automatically classify issues with Qwen3 4B Thinking (Yang et al., 2025), prompting the model to determine whether an issue reports a paper–code discrepancy or not, resulting in $232$ candidates (prompt in §F.1). All candidates are subsequently manually filtered to ensure they meet our definition of a paper-code discrepancy, yielding $59$ discrepancies.
@@
We collect reproducibility reports from the ML Reproducibility Challenge (Pineau et al., 2019; Sinha et al., 2020, 2022, 2023) which invites participants to reproduce papers from leading ML, CV, and NLP conferences (e.g., NeurIPS, ICML, ICLR, CVPR, ICCV, ACL, EMNLP) and papers from the reproducibility tracks of SIGIR and ECIR from 2020–2025. We retain only reports that reproduce a single, open-access paper with an open-source implementation, resulting in $171$ reproducibility papers. From each report, we extract mentions of paper–code discrepancies using GPT-5 resulting in $132$ candidates (prompt in §F.3). Each extracted discrepancy is manually verified to confirm adherence to our definition, of which $65$ withstand our filter.
@@
Finally, all manually filtered discrepancies are validated by Gemini 3.1 Pro and GPT-5, checking the discrepancy’s existence given the original paper, codebase, and the GitHub issue or the description extracted from the reproducibility paper. For the cases where the LLMs disagreed, we manually verified the discrepancy. During this step, the LLMs are further tasked with generating a standardized discrepancy description. We prompt the model to output 3-8 sentences and state what the paper describes, what the code implements, and where the difference is. We use these generated descriptions as the ground truth to ensure the same format and level of verbosity for paper-code discrepancies (prompt in §F.2 and §F.4). We use two state-of-the-art LLMs to perform the validation and rephrasing to prevent potential bias in the selection and phrasing (see §C.5 for an analysis). The verification step ensures adherence to the discrepancy definition, acting as a high-precision filter. By retaining only discrepancies that can be verified given the explicit evidence, we prioritize the precision and objectivity of the ground truth. In total, we obtain $92$ discrepancies from $72$ papers, where $42$ discrepancies originate from GitHub issues and $50$ from reproducibility papers.
@@
To scale the data in size and beyond the CS/AI domain, we employ synthetic data generation. Specifically, based on our GitHub crawl, we sample repositories linked to arXiv papers and those with permissive code licenses (i.e., MIT, Apache 2.0, CC-BY, BSD), as we will be redistributing their code partially. We randomly select $102$ repositories where the paper’s arXiv subject is CS (balanced over Machine Learning (ML), cs.LG, Computer Vision (CV), cs.CV, and Natural Language Processing (NLP), cs.CL), and $102$ for non-CS papers. [^2] We manually checked the repositories and removed codebases from our sample that were not suitable, such as repositories for survey papers or with very few files. Next, we prompt GPT-5 with the paper and code to generate five code diffs according to our discrepancy definition (prompt in §F.5). We then sample up to three of these changes, ensuring that they do not manipulate the same file and that the generated diff can be found by exact match in the original code. Finally, we add $13$ additional discrepancies that are labeled as paper omissions as we later find these to be most critical (see §3.2 and §5). This yields a total of $543$ discrepancies across the $204$ papers, of which $286$ discrepancies are from the CS domain, and $257$ are from other computational domains, such as Electrical Engineering and Systems Science ($114$), Physics ($51$), or Statistics ($40$).
Example data, including discrepancy type and category annotations, are shown in Table 11. Examples from the synthetic subset, including the code changes, are shown in Table 12. We release SciCoQA publicly under $C$-BY-4.0.